OpenStudy (anonymous):

x=(y+2)^2 give the domain and range and graph

OpenStudy (tkhunny):

Immediate inspection for part of it. \(x \ge 0\). Do you see why this is obvious?

OpenStudy (anonymous):

nope

OpenStudy (tkhunny):

When squaring Real Numbers, what is the result? Positive? Negative? Zero?

OpenStudy (anonymous):

whichever right?

OpenStudy (anonymous):

lol

OpenStudy (tkhunny):

(1)^2 = 1 (-1)^2 = 1 (2)^2 = 4 (-2)^2 = 4 (3)^2 = 9 (-3)^2 = 9 See any negative numbers on the right-hand side?

OpenStudy (anonymous):

ohh yeahh no there are no negatives

OpenStudy (tkhunny):

That's it. So that (y+2)^2 CANNOT be NEGATIVE. Makes sense?

OpenStudy (anonymous):

yup lol

OpenStudy (tkhunny):

This gives the Domain: \(x\ge 0\). How about the Range? What values for 'y' are acceptable?

OpenStudy (anonymous):

any. but i put in -1, 0, 1, and 2

OpenStudy (tkhunny):

You can "put in" until your paper runs out. Any value will do. This gives the Range \(-\infty<y<\infty\)

OpenStudy (anonymous):

okay, so how can i graph it. do i just plug in the numbers for y and graph it?

OpenStudy (tkhunny):

You should pick a few values around y = -2, maybe y = -4, -3, -2, -1, 0 and get a feel of it. You should not need more than that. See if you can figure out why I selected y = -2 as the central value.

OpenStudy (anonymous):

those are the numbers that chose to plug in after i saw how big the result was for the others. the central value is -2 because it equals 0

OpenStudy (tkhunny):

It is hoped that you can look at the original equation and see that this is so. y = -2 ==> x = 0. It is a good thing to notice right up front.